OpenStudy (anonymous):

The length and width of a rectangle are in the ratio 5:3. The perimeter of the rectangle is 32cm. Find the length and width.

OpenStudy (campbell_st):

use the unitary method... add the parts in the ratio... Parts = 5 + 3 and you know perimeter = 2l + 2w so 32 = 2( l + w) or 16 = l + w take the value of 16 and divide by the number of ratio parts... then length = 5 x answer width = 3 x answer

OpenStudy (anonymous):

32=5a*3a a=4 L=20 W = 12

OpenStudy (anonymous):

so the length is 10 and the width is 6?

OpenStudy (anonymous):

32 DIVIDE BY 8 = 4 so a = 4

OpenStudy (anonymous):

8 is from 5 +3

OpenStudy (campbell_st):

@bRidingHoo sight problem is length = 20 and width = 12 the perimeter is 64

OpenStudy (campbell_st):

thats correct @Sunshine447